Voir en


Accelerator Report: Advancing smoothly through recommissioning and overcoming challenges


On 16 February, the symbolic LHC key was officially handed over to the Operations team, marking the start of the hardware recommissioning phase for the LHC. Early this week, 8600 hardware tests were successfully executed, out of the 10 336 tests that have to be performed. The injection of the first protons into the LHC is scheduled for 11 March, but could be brought forward by a few days if all goes well.

Meanwhile, the commissioning of the LHC injector chain is advancing smoothly. Linac4 is consistently delivering beams to the PS Booster, which has now set up and tuned all operational beams for downstream users. The PS has also set up the initial beams required by the SPS, including a single-bunch LHC beam and a low-intensity version of the beam that is destined for the SPS North Area experiments.

On the SPS front, hardware commissioning is proceeding according to plan, though not without some challenges. The “heat run”, a critical step in the commissioning process, involves pulsing the main magnets intensively for about 12 hours, with a current close to the maximum average current the magnets can operate with. The aim of the heat run is to detect any anomalies in the cooling of the magnets. Under normal conditions, the magnets will reach their steady state temperature after 20 to 30 minutes.

Last week, only 10 minutes after the beginning of the heat run, the operators in the CERN Control Centre (CCC) received an alarm: one or more magnets in one of the SPS sextants (one sixth of the machine’s circumference) had overheated. The magnets are protected against overheating; this protection cuts the electrical current circulating through the magnets to avoid any damage. A prompt intervention by experts equipped with thermal cameras identified the problematic magnet, and an investigation revealed that rubber debris from an anti-return valve was obstructing the cooling water flow. The issue was addressed by cleaning all 110 filters in the affected sextant and flushing the circuit, ensuring that no rubber debris was left. A second heat run confirmed that all the magnets are now operating at normal temperatures.

Overall, both hardware and beam commissioning activities across the accelerator complex are progressing well. Despite some issues, typical of the annual recommissioning, nothing has jeopardised the beam delivery schedule so far.

This thermal image clearly shows a significant temperature difference between the magnet on the left (in blue) and the overheating magnet on the right (in yellow). (Image: CERN)
The water filter blocked with rubber debris. (Image: CERN)